OpenStudy (anonymous):

Which three ordered pairs complete the table to give solutions of the equation y = x – 7? A. (–3, –10), (0, –7), and (3, 4) B. (–3, –10), (0, –7), and (3, –4) C. (–3, –4), (0, –7), and (3, –4) D. (–3, 4), (0, –7), and (3, –4)

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

@leonardo0430

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

anyone

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

Try plugging all the points in the equation and see if they satisfy the equation The set/choice that has ALL the points that satisfy the equation will be the right one. For example I'll do choice A for you and you'll can do the rest. You have the equation IN choice A you have: (–3, –10), (0, –7), and (3, 4) You have the equation \(\sf y = x – 7\) plug one point from choice A : (-3, -10) \(\sf y = x – 7\\-10=-3-7\\-10=-10\) does this satisfy the equation? yes next point is (0,-7) \(\sf y = x – 7\\-7=0-7\\-7=-7\) does this satisfy the equation? yes do this with the 3rd point in choice A, to check whether A is the right one, if not, you have to keep doing this until you find the choice that has points which satisfy the equation.

OpenStudy (anonymous):

so the answer is A?

OpenStudy (anonymous):

i dont know, you have to check it. can you plug the 3rd point in A so that we can see if it satisfy the equation? note: i never like giving direct answers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

show me your solution 'coz i haven't calculated it yet

OpenStudy (anonymous):

i have no idea what to do here

OpenStudy (anonymous):

read and understand what I posted above, and tell me which part you don't understand. I'll explain it

OpenStudy (anonymous):

I think the answer is b

OpenStudy (anonymous):

are you sure? tell me why you choose B, to make sure you're not guessing (:

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

....??

OpenStudy (anonymous):

do you know how to plug points in an equation?

OpenStudy (anonymous):

because sense y=x -7 i though that it would have to be b because 3,-4 as i see all of them are - so that made me think it was be

OpenStudy (anonymous):

oh no i dont know how to plug points in equation and i only have 2 more min left till test time is up

OpenStudy (anonymous):

so sense i can find my answer ill just have to go with b

OpenStudy (anonymous):

g

OpenStudy (anonymous):

how will you pass the test or even the subject if you'll not understand the concept? guessing the answers will not help you. and no offense but openstudy has an Anti-Cheating Policy, which you can read here: \(\sf \href{http:///openstudy.com/terms-and-conditions}{Terms\ and\ Conditions}\). It clearly states there that posting test questions or quizzes are considered plagiarism/cheating, so we're not supposed help anyone who are having assessments like this.I would really like to help you understand this concept but since that's a test and this is not the right time to understand things, I'm sorry, I'm afraid I cannot help you. IF this is a homework question, I'm willing to help.

OpenStudy (anonymous):

this is homework

OpenStudy (anonymous):

yeah you stink at helping

OpenStudy (anonymous):

so goodbye

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @Warrior678 oh no i dont know how to plug points in equation and \(\sf \color{red}{ \text{i only have 2 more min left till test time is up}}\) \(\color{blue}{\text{End of Quote}}\) you just said it before, no need to defend your argument. I understand that you are offended with this, so I'm not going to say anything (: Just goodluck with you studies :)

OpenStudy (anonymous):

@Data_LG2 its just that im in k12 i dont learn nothing there so shut up

OpenStudy (anonymous):

haha im just kidding dont shut up

OpenStudy (anonymous):

:) I'm just informing you about this policy of openstudy, so that in the future no one's gonna tell you about this in a very rude way as I've seen before with other users like you. As Mark Twain said "Never let your school interfere with your education" , it's up to you whether you want to help yourself to learn or not. "You are your own worst enemy". If you don't want to educate yourself, it's your fault. It's not right to blame others just to make some alibi that you are not learning properly. It's up to you what do you want to do with your future. Institutions, teachers, peers or even your parents are only there to guide and not to do things for you because there will come a time in which you have to do things on your own without holding into someone. This is just my advice based from my own experiences. I've been in your position once.